# NeuroPhysNet: A FitzHugh-Nagumo-Based Physics-Informed Neural Network Framework for Electroencephalograph (EEG) Analysis and Motor Imagery Classification

### v1

@@ 3.4 Feature Extraction Module
The Feature Extraction Module processes the outputs of the PINN model, specifically the FHN variables $v$ (membrane potentials) and $w$ (recovery variables). Each variable is processed independently to extract node-specific temporal features, and these features are later fused for downstream tasks. The module is invoked twice: once for $v$ and once for $w$, ensuring that the extracted features capture distinct yet complementary temporal patterns inherent to these biophysically informed variables.
The inputs to the module are tensors $\mathbf{v}\in\mathbb{R}^{B\times N\times T}$ and $\mathbf{w}\in\mathbb{R}^{B\times N\times T}$, where $B$ represents the batch size, $N$ is the number of nodes (e.g., EEG channels), and $T$ is the number of time points. To align the input tensors with the convolutional layers, the module reshapes and permutes the data into $\mathbf{v}_{\text{input}},\mathbf{w}_{\text{input}}\in\mathbb{R}^{(B\cdot N)\times 1\times T}$. This restructuring ensures that each node is treated as an independent input sequence while sharing the same weights across all nodes.
@@
followed by another max-pooling operation:
$\mathbf{v}_{2}^{\text{pool}}=\text{MaxPool1D}(\mathbf{v}_{2},k_{p},s_{p}).$ (25)
After the final pooling operation, the feature maps are flattened into one-dimensional vectors:
$\mathbf{v}_{\text{flat}}=\text{Flatten}(\mathbf{v}_{2}^{\text{pool}}),$ (26)
@@
This design ensures that the module efficiently extracts complementary features from $v$ and $w$ while maintaining computational efficiency and scalability. By extracting temporal features from the biophysical variables $v$ and $w$ output by the PINN, the module not only captures local temporal dependencies but also integrates information constrained by physical laws, transforming it into enhanced network representations. These physically informed features serve a dual purpose within the network: on the one hand, they act as an augmentation mechanism, enhancing the network’s representation capability for input data, thus improving the robustness and discriminative power of features beyond physical constraints; on the other hand, the inclusion of physical information provides a solid foundation for interpretability, ensuring that the final output aligns with the fundamental principles of neuronal dynamics at the feature level.
This feature enhancement strategy is particularly well-suited for handling the complex nature of EEG data. EEG signals exhibit strong spatiotemporal correlations and nonlinear characteristics, and the variables $v$ and $w$ generated by the PINN not only result from data-driven processes but also incorporate physical constraints from the FHN model. These physical insights are transformed by the feature extraction module into high-dimensional representations that are more meaningful for downstream classification or regression tasks. This approach effectively addresses the lack of physical consistency often observed in traditional neural networks, such as MLPs or standard convolutional networks, when processing EEG data, while also mitigating performance fluctuations caused by distributional biases in the data.
Moreover, embedding physical information in the feature extraction process enhances the model’s generalization capability and provides critical support for cross-domain applications, such as BCI and medical diagnostics. By deeply leveraging the features of $v$ and $w$, this design enables the network to better capture key patterns hidden within signals, significantly improving the overall performance on specific tasks, such as motor imagery classification or brain activity recognition. This method also demonstrates the potential of combining physical modeling with data-driven approaches, paving the way for future advancements in intelligent signal processing and neuronal dynamics modeling.
@@ 4.1 Loss Function and Training
The training process was conducted over 100 epochs on a desktop computer equipped with an NVIDIA GeForce RTX 4090 GPU. The proposed NeuroPhysNet framework optimizes a composite loss function that integrates both classification and physics-based objectives, aligning model predictions with both the observed EEG data and the underlying neurodynamic principles. The total loss is defined as:
@@
For inference, the framework was deployed on a desktop computer equipped with an NVIDIA GeForce RTX 2080Ti GPU. This setup efficiently processed EEG data for MI classification tasks, leveraging the neurodynamic insights embedded in the PINN architecture. By combining data-driven classification objectives with physics-based modeling, NeuroPhysNet achieves robust and interpretable performance, addressing the challenges of generalization and reliability in medical EEG applications.
@@ 4.2.1 Dataset
The BCIC-IV-2a dataset [26] is a widely used benchmark in EEG-based MI classification research, particularly in the context of BCI development. This dataset contains EEG recordings from nine subjects performing a four-class motor imagery task. The four classes include imagined movements of the left hand, right hand, both feet, and tongue. These tasks were chosen to activate distinct cortical areas, facilitating the analysis and classification of MI-related EEG signals.
Data Collection and Setup
The EEG signals were recorded using 22 EEG electrodes placed according to the international 10-20 system, along with three electrooculography (EOG) channels to monitor eye movements. The signals were sampled at 250 Hz, providing high temporal resolution suitable for analyzing cortical dynamics during motor imagery. Each subject participated in two sessions: a training session and an evaluation session. Each session comprised six runs, with 12 trials per class, resulting in a total of 288 trials per subject.
@@
For the experiments in this study, the EEG signals were preprocessed to enhance the signal-to-noise ratio and extract relevant features. The preprocessing pipeline included band-pass filtering to isolate task-relevant frequency bands (e.g., mu and beta rhythms), artifact removal using EOG channels, and segmentation into fixed-length temporal windows aligned with the onset of motor imagery tasks. These steps ensured that the dataset was ready for input into the proposed NeuroPhysNet framework.
Significance in This Study
The BCIC-IV-2a dataset serves as a robust benchmark to evaluate the proposed NeuroPhysNet framework. Its well-annotated structure and realistic challenges, such as inter-subject variability and noise, provide a comprehensive testbed for assessing the generalization and interpretability of the physics-informed neural network architecture. By leveraging this dataset, we demonstrate the ability of NeuroPhysNet to effectively integrate biophysical constraints and extract discriminative features for motor imagery classification.
@@ 4.2.2 Baseline Models
    • Filter Bank Common Spatial Pattern (FBCSP) [2]: A classical approach that applies spatial filters across multiple frequency bands to extract task-relevant features from EEG signals.
@@
    • Minimum Distance to Mean (MDM) [28]: Classifies EEG signals by minimizing geodesic distances between covariance matrices.
    • Temporal Spectral Mapping (TSM) [19]: Combines temporal and spectral features to improve classification performance.
    • SPDNet [10]: A deep learning model specifically designed to process symmetric positive definite (SPD) matrices.
    • Tensor-CSPNet [15]: Extends CSP by incorporating tensor-based feature representations.
@@
    • ConvNet [20]: A simple convolutional neural network optimized for EEG feature extraction.
    • EEGNet [18]: A compact and efficient architecture designed specifically for brain-computer interface (BCI) applications.
    • Filter Bank Convolutional Network (FBCNet) [23]: Integrates filter banks with convolutional networks to extract multi-band EEG features.
@@
4.2.3 Fusion of $v$ and $w$ with Tensor-CSPNet Features
The fusion process combines $v$ and $w$ features with the high-dimensional tensor features produced by Tensor-CSPNet’s layers. This integration occurs at the feature level, where the outputs of Tensor-CSPNet’s feature extraction module are concatenated with the physiological The fusion process combines $v$ and $w$ features with the high-dimensional tensor features produced by Tensor-CSPNet’s layers. Tensor-CSPNet, as a state-of-the-art technique, has demonstrated outstanding performance in EEG classification tasks due to its ability to capture complex spatial and frequency characteristics of EEG signals. By leveraging advanced tensor representations, Tensor-CSPNet provides a strong foundation for feature extraction, making it one of the most effective models in this domain.
This integration occurs at the feature level, where the outputs of Tensor-CSPNet’s feature extraction module are concatenated with the physiological features. The fused feature set is then passed through additional neural layers for final classification. This approach allows the model to simultaneously leverage spatial, frequency, temporal, and physiological dynamics, enhancing its ability to discriminate between EEG classes.
@@ 4.2.3 Fusion of $v$ and $w$ with Tensor-CSPNet Features
FBCSP [2] 71.29 73.39 66.13
EEGNet [18] 69.26 66.93 60.31
ConvNet [20] 70.42 65.89 57.61
FBCNet [23] 75.48 77.16 71.53
MDM [28] 62.96 59.49 50.74
TSM [19] 68.71 63.32 49.72
SPDNet [10] 65.91 61.16 55.67
Tensor-CSPNet [15] 75.11 77.36 73.61
@@ 5 Experimental Results
Filter Bank Common Spatial Pattern (FBCSP) [2], a classical CSP-based method, achieves CV accuracies of 71.29% (T) and 73.39% (E), and a Holdout accuracy of 66.13%. While its ability to extract task-relevant spatial and frequency features is notable, its performance is surpassed by most modern methods due to its inability to model complex temporal and physiological dynamics.
For deep learning-based approaches, ConvNet [20] achieves CV accuracies of 70.42% (T) and 65.89% (E), and a Holdout accuracy of 57.61%. EEGNet [18] performs slightly worse, with CV accuracies of 69.26% (T) and 66.93% (E), and a Holdout accuracy of 60.31%. FBCNet [23], leveraging filter banks to extract multi-band features, achieves CV accuracies of 75.48% (T) and 77.16% (E), and a Holdout accuracy of 71.53%. Despite its superior frequency representation, it does not incorporate physiological dynamics, which constrains its generalization capabilities.
Among Riemannian geometry-based methods, Minimum Distance to Mean (MDM) [28] achieves CV accuracies of 62.96% (T) and 59.49% (E), with a Holdout accuracy of 50.74%. Similarly, Temporal Spectral Mapping (TSM) [21] records CV accuracies of 68.71% (T) and 63.32% (E), and a Holdout accuracy of 49.72%. SPDNet [10] shows slightly better performance than MDM, with CV accuracies of 65.91% (T) and 61.16% (E), and a Holdout accuracy of 55.67%. However, these methods struggle to handle the high-dimensional temporal dynamics inherent in EEG signals, resulting in comparatively low accuracy across all scenarios.
Tensor-CSPNet [15], which extends CSP with tensor-based feature representations, achieves CV accuracies of 75.11% (T) and 77.36% (E), and a Holdout accuracy of 73.61%. This method demonstrates competitive performance by effectively capturing spatial and frequency characteristics. However, it lacks the integration of physiological insights, limiting its ability to model the temporal dynamics of EEG signals comprehensively.
NeuroPhysNet, on the other hand, achieves the highest performance among all methods in multiple scenarios. It records CV accuracies of 76.23% (T) and 78.03% (E), and a Holdout accuracy of 74.20%. This improvement can be attributed to its innovative integration of the FHN equations. By incorporating the computed membrane potential ($v$) and recovery variable ($w$) into Tensor-CSPNet’s feature space, NeuroPhysNet enhances its ability to capture both temporal and physiological dynamics. This fusion allows it to outperform traditional CSP-based methods, Riemannian geometry-based models, and deep learning architectures, while maintaining consistent and robust generalization across diverse evaluation scenarios.
5.2 Impact of Training Data Proportion on NeuroPhysNet Performance
@@ 6 Evaluating the Robustness of NeuroPhysNet with Full and Limited Training Data
As a model based on PINNs, NeuroPhysNet inherits one of the prominent advantages of PINNs: the ability to perform well even with limited training data. This capability arises from PINNs’ integration of known physical laws, such as the FitzHugh-Nagumo equations, into the learning process of neural networks. By incorporating physical constraints, NeuroPhysNet reduces its reliance on large-scale datasets and extracts meaningful features even in data-limited scenarios, enhancing its generalization capability.
@@
These results highlight NeuroPhysNet’s superior ability to generalize and remain robust under limited data conditions. By incorporating physiological insights through the FHN equations, NeuroPhysNet captures essential temporal and physiological dynamics, allowing it to outperform Tensor-CSPNet, which primarily relies on tensor-based spatial and frequency representations without physiological integration.
@@ 6.1 Evaluating the Impact of $v$ and $w$ Features in NeuroPhysNet
To assess the individual contribution of the physiological features computed by the FHN equations, an experiment was conducted by training and testing NeuroPhysNet using only the membrane potential ($v$) and recovery variable ($w$) as input features. This experiment isolates the physiological components from other traditional EEG features to evaluate their standalone effectiveness in EEG classification tasks.
@@
Figure 1: Visualization of $v$ (membrane potential) during the Wake stage and $w$ (recovery variable) during the REM stage for selected EEG channels. The temporal dynamics of $v$ and $w$ across different channels highlight their variability and potential as features for EEG classification.
The results presented in Table 3 show that using only $v$ and $w$ features, NeuroPhysNet achieves reasonable classification performance, though it is significantly lower than the full-feature model. With the full training dataset, the model achieves a CV accuracy of 61.23% (E) and a Holdout accuracy of 59.87%. As the training data proportion decreases, performance declines but remains relatively stable. For example, at 30% training data, the CV accuracy is 50.89% (E) and 50.11% (T), and the Holdout accuracy is 49.12%. The results indicate that $v$ and $w$ features, derived from the FitzHugh-Nagumo equations, provide valuable discriminative information for EEG classification. Despite the reduced accuracy compared to the full-feature model, the performance remains relatively consistent across different data proportions. This stability demonstrates the robustness of PINNs, particularly in scenarios with limited training data. Moreover, the results emphasize the need to integrate physiological features with traditional EEG features. The full-feature NeuroPhysNet achieves significantly better accuracy by leveraging the complementary strengths of multiple feature types. Future work could explore more efficient methods for fusing these features to further enhance the model’s performance and robustness.
Figure LABEL:fig:v_w_visualization illustrates the temporal dynamics of $v$ and $w$ features for selected EEG channels during the Wake and REM stages, respectively. The variability in frequency, amplitude, phase, and noise across the channels underscores the diverse nature of these features. Such diversity reflects the underlying physiological complexity captured by the FitzHugh-Nagumo equations, which enhances the capability of NeuroPhysNet to model EEG data effectively.
In conclusion, this experiment validates the value of $v$ and $w$ features while highlighting the advantage of combining them with other EEG features. The findings reaffirm the potential of PINNs in data-limited scenarios and their effectiveness in leveraging physiological insights for EEG classification.
@@ 7 Conclusion
In this study, we explored the effectiveness of NeuroPhysNet, a model leveraging the FHN equations within the framework of PINNs, for EEG classification tasks. The incorporation of physiological features, specifically the membrane potential ($v$) and recovery variable ($w$), was analyzed to assess their standalone contributions and their role in combination with traditional EEG features.
The results demonstrate that $v$ and $w$ features, derived from the FHN equations, capture critical physiological dynamics that contribute significantly to EEG classification, even when used independently. Although their performance is lower than the full-feature model, the use of $v$ and $w$ remains robust across varying training data proportions, highlighting the stability and reliability of PINNs in data-limited scenarios. Furthermore, the integration of $v$ and $w$ with traditional EEG features significantly improves classification accuracy, underlining the complementary nature of these feature sets.
Comparisons with Tensor-CSPNet reveal that NeuroPhysNet achieves superior accuracy and robustness, particularly when training data is limited. The ability of NeuroPhysNet to generalize effectively stems from its integration of physiological insights, which allows it to model complex temporal and spatial dynamics inherent in EEG signals. Additionally, the visualization of $v$ and $w$ features illustrates their diverse temporal patterns across different sleep stages, reinforcing their value as discriminative features.
Overall, this work highlights the potential of PINNs, such as NeuroPhysNet, in addressing challenges in EEG classification, especially in scenarios with constrained data availability. By leveraging physiological equations and integrating them with conventional methods, NeuroPhysNet sets a benchmark for combining domain knowledge with data-driven learning. Future research could focus on optimizing feature fusion strategies, exploring other physiological models, and extending the applicability of NeuroPhysNet to broader EEG datasets and tasks.
@@
7.0.1 Acknowledgements
The authors gratefully acknowledge the computational resources provided by the University of Technology Sydney. Additionally, AI-assisted technology (ChatGPT-4) was utilized for language editing and grammatical refinement.
[…]
The authors declare no conflicts of interest.

### v2

@@ 3.4 Feature Extraction Module
The Feature Extraction Module processes the outputs of the PINN model, specifically the FHN variables $v$ (membrane potentials) and $w$ (recovery variables). The inputs to the module are tensors $\mathbf{v}\in\mathbb{R}^{B\times N\times T}$ and $\mathbf{w}\in\mathbb{R}^{B\times N\times T}$, where $B$ represents the batch size, $N$ is the number of nodes (e.g., EEG channels), and $T$ is the number of time points. To align the input tensors with the convolutional layers, the module reshapes and permutes the data into $\mathbf{v}_{\text{input}},\mathbf{w}_{\text{input}}\in\mathbb{R}^{(B\cdot N)\times 1\times T}$.
@@
followed by another max-pooling operation and the feature maps are flattened into one-dimensional vectors:
$\mathbf{v}_{\text{flat}}=\text{Flatten}(\text{MaxPool1D}(\mathbf{v}_{2},k_{p},s_{p})),$ (25)
@@ 4.1 Loss Function and Training
The training process was conducted over 100 epochs on a computer equipped with an NVIDIA GeForce RTX 4090 GPU. The total loss is defined as:
@@ 4.2.1 Dataset
The BCIC-IV-2a dataset [tangermann2012review] is a widely used benchmark in EEG-based MI classification research, particularly in the context of BCI development. This dataset contains EEG recordings from nine subjects performing a four-class motor imagery task. The four classes include imagined movements of the left hand, right hand, both feet, and tongue.
@@
For the experiments in this study, the EEG signals were preprocessed to enhance the signal-to-noise ratio and extract relevant features. The preprocessing pipeline included band-pass filtering to isolate task-relevant frequency bands (e.g., mu and beta rhythms), artifact removal using EOG channels, and segmentation into fixed-length temporal windows aligned with the onset of motor imagery tasks.
@@ 4.2.2 Baseline Models
    • Filter Bank Common Spatial Pattern (FBCSP) [ang2008filter]: A classical approach that applies spatial filters across multiple frequency bands to extract task-relevant features from EEG signals.
@@
    • Minimum Distance to Mean (MDM) [tevet2022human]: Classifies EEG signals by minimizing geodesic distances between covariance matrices.
    • Temporal Spectral Mapping (TSM) [lin1811temporal]: Combines temporal and spectral features to improve classification performance.
    • SPDNet [huang2017riemannian]: A deep learning model specifically designed to process symmetric positive definite (SPD) matrices.
    • Tensor-CSPNet [ju2022tensor]: Extends CSP by incorporating tensor-based feature representations.
@@
    • ConvNet [liu2023novel]: A simple convolutional neural network optimized for EEG feature extraction.
    • EEGNet [lawhern2018eegnet]: A compact and efficient architecture designed specifically for brain-computer interface (BCI) applications.
    • Filter Bank Convolutional Network (FBCNet) [mane2021fbcnet]: Integrates filter banks with convolutional networks to extract multi-band EEG features.
@@
FBCSP [ang2008filter] 71.29 73.39 66.13
EEGNet [lawhern2018eegnet] 69.26 66.93 60.31
ConvNet [liu2023novel] 70.42 65.89 57.61
FBCNet [mane2021fbcnet] 75.48 77.16 71.53
MDM [tevet2022human] 62.96 59.49 50.74
TSM [lin1811temporal] 68.71 63.32 49.72
SPDNet [huang2017riemannian] 65.91 61.16 55.67
Tensor-CSPNet [ju2022tensor] 75.11 77.36 73.61
@@ 5 Experimental Results
Filter Bank Common Spatial Pattern (FBCSP) [ang2008filter], a classical CSP-based method, achieves CV accuracies of 71.29% (T) and 73.39% (E), and a Holdout accuracy of 66.13%. While its ability to extract task-relevant spatial and frequency features is notable, its performance is surpassed by most modern methods due to its inability to model complex temporal and physiological dynamics.
For deep learning-based approaches, ConvNet [liu2023novel] achieves CV accuracies of 70.42% (T) and 65.89% (E), and a Holdout accuracy of 57.61%. EEGNet [lawhern2018eegnet] performs slightly worse, with CV accuracies of 69.26% (T) and 66.93% (E), and a Holdout accuracy of 60.31%. FBCNet [mane2021fbcnet], leveraging filter banks to extract multi-band features, achieves CV accuracies of 75.48% (T) and 77.16% (E), and a Holdout accuracy of 71.53%. Despite its superior frequency representation, it does not incorporate physiological dynamics, which constrains its generalization capabilities.
Among Riemannian geometry-based methods, Minimum Distance to Mean (MDM) [tevet2022human] achieves CV accuracies of 62.96% (T) and 59.49% (E), with a Holdout accuracy of 50.74%. Similarly, Temporal Spectral Mapping (TSM) [lotte2018review] records CV accuracies of 68.71% (T) and 63.32% (E), and a Holdout accuracy of 49.72%. SPDNet [huang2017riemannian] shows slightly better performance than MDM, with CV accuracies of 65.91% (T) and 61.16% (E), and a Holdout accuracy of 55.67%. However, these methods struggle to handle the high-dimensional temporal dynamics inherent in EEG signals, resulting in comparatively low accuracy across all scenarios.
Tensor-CSPNet [ju2022tensor], which extends CSP with tensor-based feature representations, achieves CV accuracies of 75.11% (T) and 77.36% (E), and a Holdout accuracy of 73.61%.
NeuroPhysNet, on the other hand, achieves the highest performance among all methods in multiple scenarios. It records CV accuracies of 76.23% (T) and 78.03% (E), and a Holdout accuracy of 74.20%.
@@ 6.1 Evaluating the Impact of $v$ and $w$ Features in NeuroPhysNet
To assess the individual contribution of the physiological features computed by the FHN equations, an experiment was conducted by training and testing NeuroPhysNet using only the membrane potential ($v$) and recovery variable ($w$) as input features, as shown in Figure 1.
Figure 1: Visualization of $v$ (membrane potential) during the Wake stage and $w$ (recovery variable) during the REM stage for selected EEG channels.
@@
The results presented in Table 3 show that using only $v$ and $w$ features, NeuroPhysNet achieves reasonable classification performance, though it is significantly lower than the full-feature model. With the full training dataset, the model achieves a CV accuracy of 61.23% (E) and a Holdout accuracy of 59.87%. As the training data proportion decreases, performance declines but remains relatively stable. For example, at 30% training data, the CV accuracy is 50.89% (E) and 50.11% (T), and the Holdout accuracy is 49.12%. The results indicate that $v$ and $w$ features, derived from the FitzHugh-Nagumo equations, provide valuable discriminative information for EEG classification. Despite the reduced accuracy compared to the full-feature model, the performance remains relatively consistent across different data proportions.
@@ 7 Conclusion
In this study, we evaluated NeuroPhysNet—a PINN‐based EEG classifier that integrates FHN dynamics—by examining the standalone and combined contributions of the membrane potential $v$ and recovery variable $w$. These physiological features proved robust in data‐limited scenarios and, when fused with conventional EEG features, yielded significant accuracy gains. Compared to previous networks, NeuroPhysNet demonstrated superior accuracy and generalization. Visualization of $v$ and $w$ across sleep stages further confirmed their discriminative power. Our results highlight the promise of PINNs for EEG classification under constrained data conditions and point toward future work on refined fusion strategies, alternative physiological models, and broader EEG applications.